# Robustness of topological phases on aperiodic lattices

### v1

@@ 4.2. Stacked topological phases are weak
where $\mu_{\Lambda}$ and $\mu_{L}$ are the images of $\mu$ under the coordinate projections $\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ and $\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$. This finishes the proof. ∎
Corollary 4.7 allows us to define a *-homomorphism between groupoid C*-algebras. Since $\mathcal{G}_{L}$ is an étale groupoid, its C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{L})$ has a unit, given by the constant function $1_{\Omega_{0}^{L}}$ on the unit space. Define
@@
Then $\varphi^{\mathrm{Gpd}}$ gives a *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$ under the isomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\simeq\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}^{*}(\mathcal{G}_{L})$ in Corollary 4.7. We write $\varphi^{\mathrm{Gpd}}_{N}\colon\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$ for its entrywise extension to matrix algebras.
@@
So $\varphi^{\mathrm{Roe}}(T)$ is locally compact or controlled iff $T$ is locally compact or controlled (cf. Example 3.15). Thus $\varphi^{\mathrm{Roe}}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$.
@@
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))$
for all $k$.
@@
+:$\ell_{+}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{n-1}\times\mathbb{R}_{\geq 0}),\quad\ell_{-}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{n-1}\times\mathbb{R}_{<0}),$
@@
Let $\xi^{\mathrm{Gpd}}_{\omega\times\ell,N}$ be the spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\otimes\mathbb{M}_{N}(\mathbb{R})$ defined in (3.10). Then its induced map
$\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))\to\mathbb{Z}$
@@
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{m+n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{m+n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})).$
@@
hold for all $x,y\in\omega$ and $a,b\in\ell$. Therefore the maps $\varphi^{\mathrm{Roe}}\circ\pi_{\omega}^{N}$ coincides with $\pi_{\omega\times\ell}^{N}\circ\varphi^{\mathrm{Gpd}}$ on all elements of the form $f\otimes S$, and hence for $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. That is, the diagram above commutes.
The commutative diagram above gives a commutative diagram in K-theory. This, together with the the bottom left triangle of (4.2), gives the following commutative diagram:
${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))}$${\mathbb{Z}}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))}$${\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))}$$\scriptstyle{\varphi^{\mathrm{Gpd}}_{*}}$$\scriptstyle{\pi_{\omega}}$$\scriptstyle{\pi_{\omega\times\ell}}$$\scriptstyle{\xi^{\mathrm{Gpd}}_{\omega\times\ell}}$$\scriptstyle{\varphi^{\mathrm{Roe}}_{*}}$$\scriptstyle{\xi^{\mathrm{Roe}}_{\omega\times\ell}}$
@@
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)),$
which vanishes by Lemma 4.10. Therefore $\xi^{\mathrm{Gpd}}_{\omega\times\ell}\circ\varphi^{\mathrm{Gpd}}_{*}$ must be the zero map. ∎
@@
Let $\Lambda=\mathbb{Z}^{d}$ and $L=\mathbb{Z}$. Then $\mathcal{G}_{\Lambda}\simeq\mathbb{Z}^{d}$ and $\mathcal{G}_{L}\simeq\mathbb{Z}$ as topological groupoids. The unit spaces $\Omega_{0}^{\Lambda}$ and $\Omega_{0}^{\Lambda\times L}$ are both singletons. So there are unique localised regular representations $\pi\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ and $\mathrm{C}^{*}(\mathbb{Z}^{d+1})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d+1})$. Let $N=1$. The *-homomorphism $\varphi^{\mathrm{Gpd}}_{1}\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ maps an element in $\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ to its restriction to the first $d$-coordinates. If we identify $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with $\mathrm{C}(\mathbb{T}^{d})$ via Fourier transform, then the map $\varphi^{\mathrm{Gpd}}$ sends $f\in\mathrm{C}(\mathbb{T}^{d})$ to $f\otimes 1\in\mathrm{C}(\mathbb{T}^{d})\otimes\mathrm{C}(\mathbb{T})=\mathrm{C}(\mathbb{T}^{d+1})$. The map $\varphi^{\mathrm{Gpd}}$ is induced by the injective group homomorphism
@@
hence its image belongs to the kernel of the map
$\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))$
by [Ewert-Meyer:Coarse_geometry]*Proposition 10.

### v2

@@ 4.2. Stacked topological phases are weak
where $\mu_{\Lambda}$ and $\mu_{L}$ are the images of $\mu$ under the coordinate projections $\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}^{p}$ and $\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}^{q}$. This finishes the proof. ∎
Corollary˜4.6 allows us to define a *-homomorphism between groupoid C*-algebras. Since $\mathcal{G}_{L}$ is an étale groupoid with compact unit space, its C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{L})$ has a unit, given by the constant function $1_{\Omega_{0}^{L}}$ on the unit space. Define
@@
Then $\varphi^{\mathrm{Gpd}}$ gives a *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})$ under the isomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\simeq\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathrm{C}^{*}(\mathcal{G}_{L})$ in Corollary˜4.6. We write $\varphi^{\mathrm{Gpd}}_{N}\colon\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$ for its entrywise extension to matrix algebras.
@@
So $\varphi^{\mathrm{Roe}}(T)$ is locally compact or controlled if and only if $T$ is locally compact or controlled (cf. Example˜3.18). Thus $\varphi^{\mathrm{Roe}}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)$.
@@
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))$
for all $n$.
@@
$\ell_{+}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{q-1}\times\mathbb{R}_{\geq 0}),\quad\ell_{-}\mathrel{\vcentcolon=}\ell\cap(\mathbb{R}^{q-1}\times\mathbb{R}_{<0}),$
@@
Let $\xi^{\mathrm{Gpd}}_{\omega\times\ell,N}$ be the spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})\otimes\mathbb{M}_{N}(\mathbb{R})$ defined in (3.13). Then for each $n$, its induced map
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))\to\operatorname{KK}(\mathrm{C}\ell_{n,p+q},\mathbb{R})$
@@
$\varphi^{\mathrm{Gpd}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\operatorname{KK}(\mathrm{C}\ell_{n},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L})).$
@@
holds for all $x,y\in\omega$ and $a,b\in\ell$. Therefore the maps $\varphi^{\mathrm{Roe}}\circ\pi_{\omega}^{N}$ coincides with $\pi_{\omega\times\ell}^{N}\circ\varphi^{\mathrm{Gpd}}$ on all elements of the form $f\otimes S$, and hence for $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. That is, the diagram above commutes.
The commutative diagram above gives a commutative diagram in K-theory. This, together with the bottom left triangle of (4.2), gives the following commutative diagram:
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathcal{G}_{\Lambda\times L}))$$\operatorname{KK}(\mathrm{C}\ell_{n,p+q},\mathbb{R})$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell))$$\scriptstyle{\varphi^{\mathrm{Gpd}}_{*}}$$\scriptstyle{\pi_{\omega}}$$\scriptstyle{\pi_{\omega\times\ell}}$$\scriptstyle{\xi^{\mathrm{Gpd}}_{\omega\times\ell}}$$\scriptstyle{\varphi^{\mathrm{Roe}}_{*}}$$\scriptstyle{\xi^{\mathrm{Roe}}_{\omega\times\ell}}$
@@
$\varphi^{\mathrm{Roe}}_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega\times\ell)),$
which vanishes by Lemma˜4.9. Therefore $\xi^{\mathrm{Gpd}}_{\omega\times\ell}\circ\varphi^{\mathrm{Gpd}}_{*}$ must be the zero map. ∎
@@
Let $\Lambda=\mathbb{Z}^{d}$ and $L=\mathbb{Z}$. Then $\mathcal{G}_{\Lambda}\simeq\mathbb{Z}^{d}$ and $\mathcal{G}_{L}\simeq\mathbb{Z}$ as topological groupoids. The unit spaces $\Omega_{0}^{\Lambda}$ and $\Omega_{0}^{\Lambda\times L}$ are both singletons. So there are unique localised regular representations $\pi\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ and $\mathrm{C}^{*}(\mathbb{Z}^{d+1})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d+1})$. Let $N=1$. The *-homomorphism $\varphi^{\mathrm{Gpd}}_{1}\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\to\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ maps an element in $\mathrm{C}^{*}(\mathbb{Z}^{d+1})$ to its restriction to the first $d$-coordinates. Identify $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with $\mathrm{C}(\mathbb{T}^{d})$ via Fourier transform, then the map $\varphi^{\mathrm{Gpd}}$ sends $f\in\mathrm{C}(\mathbb{T}^{d})$ to $f\otimes 1\in\mathrm{C}(\mathbb{T}^{d})\otimes\mathrm{C}(\mathbb{T})=\mathrm{C}(\mathbb{T}^{d+1})$. The map $\varphi^{\mathrm{Gpd}}$ is induced by the injective group homomorphism
@@
and its image belongs to the kernel of the map
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})),$
which yields [Ewert-Meyer:Coarse_geometry]*Proposition 10.
Acknowledgement
The author is grateful to the anonymous referee for various illuminating suggestions, which significantly improved the results of this work. The author also thanks Chris Bourne, Matthias Frerichs, Matthias Ludewig and Guo Chuan Thiang for helpful discussions, and his PhD supervisor Bram Mesland for heartfelt encouragement and a careful proofreading, helping the author to eliminate various errors in an earlier version of this work. The work is part of NWO project 613.009.142 “Noncommutative index theory of discrete dynamical systems”.
Author declarations
Conflict of interest
The author has no conflict of interest to disclose.
Data availability
Data sharing is not applicable in this study as no new data were created or analysed.